Fermi National Laboratory


Accelerator Update for October 17, 2005 - October 19, 2005

Monday October 17
The day shift began with Operations monitoring store 4449 and a Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with NuMI off for chiller work, and with the MiniBooNE and the Meson MCenter experiments taking beam.

Operations began preparing the Recycler at 8:24 AM to stash antiprotons.

At 9:19 AM, Operations reported another vacuum burst in the Pbar A150 transfer line. Two Tevatron (TeV) sector F0 Lambertson vacuum pumps tripped off. This caused the beam to move a bit, but there was no TeV abort this time.

The Recycler stash ended. Pbar and MCenter resumed taking beam by 9:54 AM.

A power supply expert reported at 11:25 AM that he had completed repairs on the Meson MTest power supply (MT5Q2), which had sprung a LCW leak.

A Pbar expert began a target spot size study at 12:20 PM. The study lasted for 2.5 hours.

At 2:40 PM, Operations reported a Recycler vacuum burst. The vacuum pumps that tripped off reset okay. The stash suffered from an emittance blowup, but no antiprotons were lost.

Recycler began preparing at 3:15 PM to stash. The stash ended and Pbar resumed stacking at 4:21 PM.

The Pbar expert resumed his studies at 4:32 PM.

MCenter resumed taking beam at 4:56 PM.

The NuMI work ended at 5:54 PM. Operations began reestablishing beam to the experiment.

A TeV expert began an end-of-store study at 8:04 PM.

The Recycler began preparing around 9:30 PM to stash.

Operations terminated store 4449 at 9:35 PM. Operations prepared CDF and D0 for a collision hall access.

At 10:28 PM, a vacuum burst in TeV sector F0 worked its way into the Main Injector (MI) P150 transfer line. Operators reset vacuum pumps.

CDF dropped their interlocks at 11:08 PM.

Tuesday October 18
The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operators completed interlocking CDF at 12:08 PM.

Operations put the TeV through a dry squeeze at 12:31 AM.

At 12:53 AM, a Pbar Debuncher pulsed magnet power supply (D:PMAGV) tripped on a ground fault indication. Operators responded to AP0 and reset the supply.

Operations put the TeV into shot setup at 12:55 AM.

At 1:08 AM, PMAGV tripped off again. Operators reset the supply, but it tripped off again. Operations called in experts.

Operations established Recycler only store 4451 at 3:18 AM with an initial luminosity of 134.8E30.

The Pbar expert reported at 3:29 AM that the PMAG power supply was okay. He suspected the problem was in the PMAG module, which would require a tunnel access. The Run Coordinator wanted Operations to start calling people in at 6 AM to begin tunnel investigation and repairs.

Operations halted NuMI beam around 7:30 AM. NuMI reported that they would be continuing their Horn chiller work today.

At 10:05 AM, an apparent site wide power glitch caused store 4451 to abort. There was no quench. Almost all the RF throughout the complex tripped off. The stash was also lost.

The High Voltage expert reported at 10:13 AM that the power glitch came from off site on the ComEd grid.

Operations prepared the TeV, MI, and Booster for access at 10:37 AM. Vacuum experts will begin the vacuum repairs in F sector. Other techs and experts will begin work everywhere else.

Pbar experts began replacing PMAG at 1:01 PM.

The Booster work finished and experts began a 400 MeV study at 2:51 PM.

Both CDF and D0 were out of access by 4:05 PM.

MI came out of access at 5:15 PM.

TeV came out of access at 6:25 PM. The vacuum experts reported that they could not find a leak in F sector.

Operations established beam to the Booster dump at 8:02 PM.

Operations put the TeV through a dry squeeze at 10:32 PM.

Operations established beam to NuMI at 10:53 PM.

A Recycler expert began a study period around 11 PM.

Operations established beam to MCenter at 11:07 PM.

Wednesday October 19
The midnight shift began with NuMI and MCenter taking beam.

Pbar experts reported at 12:08 AM that they had successfully pulsed PMAG. They began replacing the shielding blocks.

Pbar came out of access at 1:17 AM.

A Recycler expert began a study period around 1:20 AM.

Pbar began stacking at 1:41 AM.

At 3:56 AM, a Linac RF station (LRF3) tripped off for the first of five times in about three hours. It reset each time.

At 7:15 AM, Operations reported a MI P1 transfer line vacuum burst. Operations had it under control by 7:19 AM.

The Plans for Wednesday
The plans for today are to stack and stash, and hopefully put in a store later this evening. Until then, the Run Coordinator will allow TeV experts to conduct studies.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/19/2005   email Fermilab